OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 16 and 18 and a vertex at (17, -2)

OpenStudy (campbell_st):

well the equation is y =a(x -16)(x -18) substitute the point to find the value of a when you know a, just distribute

OpenStudy (anonymous):

Lolz beat me to it! @campbell_st

hero (hero):

Or if you don't want to do all that work, just input the vertex (17,-2) and the point (18,0) into this equation y = a(x - h)^2 + k as follows: 0 = a(18 - 17)^2 - 2 Then solve for a